# Exploring the Frontiers of Softmax: Provable Optimization, Applications in Diffusion Model, and Beyond

### v1

@@
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Chenyang Li lchenyang550@gmail.com. Fuzhou University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.
@@ 1 Introduction
Large Language Models (LLMs) like GPT4 [1] from OpenAI and Claude 3 [5] from Anthropic have widely and profoundly changed the world. Some researchers believe they split human history into two parts, the Pre-LLM Era and the LLM Era. The LLMs have been widely used in human activities, such as education [79], law [132], finance [89], bio-informatics [144], coding [72], and even top AI conference reviews such as ICML, ICLR, and NeurIPS [80]. To make LLMs successful, one of the cores of LLMs is the Transformer model architecture [146], which has many advantages, including faster-parallelized inference rather than sequential inference like RNN [67]; being easy to scale up the model capacity to support the scaling laws in neural language models [77], i.e. since the input and output dimension of each Transformer blocks is the same, we can stack an arbitrary number of layers as we want. The kernel design of the Transformer block is self-attention layers, where each block has many attention heads and each head has its three important private parameter matrices for key, query, and value operation. Many papers believe that the self-attention operation is the critical reason for emergent ability [150], including in-context learning [100, 113] and compositional ability to solve complex task [42, 86]. The Transformer is so successful and has been widely certified that this architecture can be adopted in many other modalities such as tabular data, image/video generation, e.g. the video diffusion model SORA [103] from OpenAI using Transformer [107] as its backbone.
()(*)~{}( ∗ ) What is the learning mechanism that makes softmax so powerful?
To demystify the black box, in this paper, we analyze the Gradient Descent (GD) training dynamics for two-layer Neural Networks (NN) with softmax activation function for multi-dimensional regression, i.e., $F(W,x,a)\in\mathbb{R}^{d}$ and
$F(W,x,a)_{\ell}:={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}~{}~{}~{}~{}\forall\ell\in\{1,\dots,d\},$
where $m$ is number of hidden neurons, $\exp(\cdot)$ is element-wise exponential function, $a_{\ell},W$ are the first and second layer weights respectively and $x$ is the input data. Note that, the self-attention could be written as $F(W^{K}X,W^{Q}X,W^{V}X)\in\mathbb{R}^{d\times n^{\prime}}$, where $W^{K},W^{Q},W^{V}\in\mathbb{R}^{d\times d}$ denotes key, query, and value matrix and $X\in\mathbb{R}^{d\times n^{\prime}}$ is a sequence of $n^{\prime}$ tokens. Thus, studying the two-layer softmax network is the prerequisite to understanding self-attention. See more discussion in Section 7.
@@
Table 1: Comparing hidden neuron number $m$ in two-layer neural networks and training steps $\widehat{T}$ are required under different activation functions to guarantee that, for any $\epsilon>0$, with probability at least $0.99$, the training loss is smaller or equal to $\epsilon$. Here, $n$ is the number of training samples and $\lambda$ is the smallest eigenvalue for the matrix of neural tangent kernel, where $n>1$ and $\lambda<1$. We can see that the two-layer NN with softmax activation function requires almost the same number of neurons and training steps to converge as that with ReLU or exponential activation functions. More details: Theorem 3.6 in [94] for ReLU; Theorem 1.1 in [53] for $\exp$; Corollary 4.3 in our paper for softmax.
@@
To answer our ()(*)( ∗ ) question above, we adopt the powerful NTK analysis paradigm in this work. Our analysis shows that, because of the normalization effect of the denominator, the Neural Tangent Kernel induced by the softmax has a good perturbation property (Lemma 5.1), which means the loss landscape of softmax version has a large convex region. Thus, the softmax NN requires almost the same number of neurons and training steps to fit the data and converge as ReLU or exponential NN, which is illustrated in Table 1 clearly (Theorem 4.2). To demonstrate the broad applicability of our theoretical findings, we apply our analysis in a practical case study to show the generalization ability of softmax NN, where the task is learning score estimation functions in diffusion models with noisy labels, a promising approach for generative modeling, as we can smartly transfer it to a multi-dimensional regression task (Theorem 6.5). Thus, we show that gradient-based algorithms can learn the score function with a provable accuracy.
@@
  • Softmax NTK: We build up the first NTK analysis framework for two-layer NN with softmax activation function. Furthermore, our multi-dimensional regression setting is more general than previous work [94, 53] (ReLU and $\exp$) and can be degenerated to the linear regression setting.
  • Diffusion Models Case Study: We apply our results in learning score estimation functions in diffusion models with noisy labels to verify our analysis effectiveness.
2 Related Work
Softmax and Attention in LLMs.
@@ 2 Related Work
Neural Tangent Kernel.
Diffusion Model.
Score-based generative diffusion models can generate high-quality image samples comparable to GANs which requires adversarial optimization [60, 129, 76]. Based on the U-Net [114], stable diffusion can successfully generate business-used images. Based on the softmax-based self-attention [107], OpenAI released a video diffusion model, SORA [103], with a surprising performance. Another line of work studying how to train the diffusion models to have a better theoretical guarantee [123, 124, 127, 125, 122, 83, 75, 121, 81, 31, 23, 27, 118, 154, 14, 48, 22, 49, 147, 28]. In this work, we adapt our analysis in diffusion models.
@@ 3 Preliminary
[…]
As we have multiple index, to avoid confusion, we usually use $i,j\in[n]$ to index the training data, $\ell\in[d]$ to index the output dimension, $r\in[m]$ to index neuron number.
Models. We consider a two-layer softmax neural network. The hidden layer has $m$ neurons, and we use the softmax function as the activation function, $F(W,\cdot,a):\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ and
$F(W,x,a)_{\ell}:={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}~{}~{}~{}~{}\forall\ell\in[d_{2}],$ (1)
where $\exp(\cdot)$ is element-wise exponential function. We use $m$ as a normalization factor. Note that we can reduce the $d_{2}$ to $1$ for the linear regression setting. To simplify the proof we let $d_{1}=d_{2}$. Note that our proof can generalize to different $d_{1},d_{2}$ easily.
@@
Data. We have $n$ training data points ${\cal D}_{n}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$. [^1] Our analysis can extend to $x_{i}\in\mathbb{R}^{d_{1}}$ and $y_{i}\in\mathbb{R}^{d_{2}}$ easily. We denote $X=[x_{1},\dots,x_{n}]\in\mathbb{R}^{d\times n}$ and $Y=[y_{1},\dots,y_{n}]\in\mathbb{R}^{d\times n}$. We assume that $\|x_{i}\|_{2}\leq 1$ and $\|y_{i}\|_{2}\leq 1$, $\forall i\in[n]$. We have the softmax function $\mathcal{S}\in\mathbb{R}^{m\times n}$, where $\mathcal{S}_{i}\in\mathbb{R}^{m}$ denotes $\langle\exp(W^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(W^{\top}x_{i})$ and $\mathcal{S}_{i,r}\in\mathbb{R}$ denotes $\langle\exp(W^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(w_{r}^{\top}x_{i})$, $\forall r\in[m],\forall i\in[n]$. For simplicity, we denote ${\alpha}_{i}$ as $\langle{\bf 1}_{m},\exp({W}^{\top}x_{i})\rangle$, $\exp_{i}$ as $\exp(W^{\top}x_{i})$ and $\exp_{i,r}$ as $\exp(w_{r}^{\top}x_{i})$, $\forall r\in[m],\forall i\in[n]$, when the context is clear.
Gradient Descent. We use $e_{r}$ to denote a vector where the $r$-th coordinate is $1$ and everywhere else is 00. $\forall r\in[m],\forall\ell\in[d]$, we have $\frac{\partial F(W,x,a)_{\ell}}{\partial w_{r}}\in\mathbb{R}^{d}$ can be written as
$\frac{\partial F(W,x,a)_{\ell}}{\partial w_{r}}=$ $~{}+{{}{}{}m}\langle a_{\ell}\circ e_{r},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}x$
$~{}-{{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-2}\cdot\langle\exp(W^{\top}x),e_{r}\circ{\bf 1}_{m}\rangle x$
@@
We use $W(\tau)$ to denote the weights of the first layer on the timestamp $\tau$ and similar for $\mathcal{S}(\tau)$ and $F(\tau)$ when the context is clear. Now, we introduce some necessary definition used.
@@
$\Delta w_{r}(\tau):={{}{}{}m}\sum_{i=1}^{n}\sum_{\ell=1}^{d}(F_{\ell,i}(\tau)-y_{\ell,i})\cdot\Big{(}\langle a_{\ell}\circ e_{r},\mathcal{S}_{i}(\tau)\rangle-\langle a_{\ell},\mathcal{S}_{i}(\tau)\rangle\cdot\langle\mathcal{S}_{i}(\tau),e_{r}\circ{\bf 1}_{m}\rangle\Big{)}\cdot x_{i}$
where $\mathcal{S}_{i}(\tau)=\langle\exp(W(\tau)^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(W(\tau)^{\top}x_{i})\in\mathbb{R}^{m}$.
Note that we can simplify the gradient calculation by the fact $1=\langle{\bf 1}_{m},\mathcal{S}_{i}(\tau)\rangle$. Thus, we have the following claim.
@@ 3.1 Neural Tangent Kernel
$[H_{\ell_{1},\ell_{2}}]_{i,j}(W):=\frac{1}{m}x_{i}^{\top}x_{j}\sum_{r=1}^{m}\langle v_{\ell_{1},r},\mathcal{S}_{i}\rangle\cdot{{}{}{}m}\mathcal{S}_{i,r}\cdot\langle v_{\ell_{2},r},\mathcal{S}_{j}\rangle\cdot{{}{}{}m}\mathcal{S}_{j,r}.$
@@
Note that $H^{*}$ is a positive semi-definite matrix, and we denote its minimum eigenvalue as $\lambda:=\lambda_{\min}(H^{*})$.
@@ 4 Main Results
Let $C>10$ denote a sufficiently large constant. We define parameter $B$ as follows
$B:=\max\{C\sigma\sqrt{\log(nd/\delta)},~{}1\}.$
@@
Let $\lambda=\lambda_{\min}(H^{*})>0$, $m=\Omega(\lambda^{-2}n^{2}d^{2}\exp(18B)\log^{2}(nd/\delta))$, $\eta=0.1\lambda/(mn^{2}d^{2}\exp(16B))$ and $\widehat{T}=\Omega((m\eta\lambda)^{-1}\log(nd/\epsilon))={{}{}{}\Omega(\lambda^{-2}n^{2}d^{2}\exp(16B)\cdot\log(nd/\epsilon))}$. For any $\epsilon,\delta\in(0,0.1)$, after $\widehat{T}$ iterations, with probability at least $1-\delta$, we have
$\|F(\widehat{T})-Y\|_{F}^{2}\leq\epsilon.$
@@
Consider the 1-dimension linear regression setting, i.e., $d_{1}=d$ and $d_{2}=1$. Let $\lambda=\lambda_{\min}(H^{*})>0$, $m=\Omega(\lambda^{-2}n^{2}\exp(18B)\log^{2}(n/\delta))$, $\eta=0.1\lambda/(mn^{2}\exp(16B))$ and $\widehat{T}=\Omega((m\eta\lambda)^{-1}\log(n/\epsilon))={{}{}{}\Omega(\lambda^{-2}n^{2}\exp(16B)\cdot\log(n/\epsilon))}$. For any $\epsilon,\delta\in(0,0.1)$, after $\widehat{T}$ iterations, with probability at least $1-\delta$, we have
$\|F(\widehat{T})-Y\|_{2}^{2}\leq\epsilon.$

### v2

@@
Yang Cao and Yingyu Liang and Zhenmei Shi and Zhao Song Wyoming Seminary.The University of Hong Kong.University of Wisconsin-Madison.magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.
@@ 1 Introduction
Large Language Models (LLMs) like GPT4 [1] from OpenAI and Claude 3 [10] from Anthropic have widely and profoundly changed the world. Some researchers believe they split human history into two parts: the Pre-LLM Era and the LLM Era. The LLMs have been widely used in human activities, such as education [75], law [141], finance [83], bio-informatics [142], coding [66], and even top AI conference reviews such as ICML, ICLR, NeurIPS, and AISTATS [86]. To make LLMs successful, one of the cores of LLMs is the Transformer model architecture [145], which has many advantages, including faster-parallelized inference rather than sequential inference like RNN [64]; being easy to scale up the model capacity to support the scaling laws in neural language models [74], i.e. since the input and output dimension of each Transformer blocks is the same, we can stack an arbitrary number of layers as we want. The kernel design of the Transformer block is self-attention layers, where each block has many attention heads and each head has its three important private parameter matrices for key, query, and value operation. Many papers believe that the self-attention operation is the critical reason for emergent ability [149], including in-context learning [100, 112] and compositional ability to solve complex task [46, 89]. The Transformer is so successful and has been widely certified that this architecture can be adopted in many other modalities such as tabular data, image/video generation, e.g., the video diffusion model SORA [102] using Transformer [107] as its backbone.
Table 1: Comparing hidden neuron number $m$ in two-layer neural networks and training steps $\widehat{T}$ are required under different activation functions to guarantee that, for any $\epsilon>0$, with probability at least $0.99$, the training loss is smaller or equal to $\epsilon$. Here, $n$ is the number of training samples, and $\lambda$ is the smallest eigenvalue for the matrix of the neural tangent kernel, where $n>1$ and $\lambda<1$. We can see that the two-layer NN with softmax activation function requires almost the same number of neurons and training steps to converge as that with ReLU or exponential activation functions. More details: Theorem 3.6 in [97] for ReLU; Theorem 1.1 in [48] for $\exp$; Corollary 4.3 in our paper for softmax.
@@
()(*)~ What is the learning mechanism that makes softmax so powerful?
To demystify the black box, in this paper, we analyze the Gradient Descent (GD) training dynamics for two-layer Neural Networks (NN) with softmax activation function for multi-dimensional regression, i.e., $F(W,x,a)\in\mathbb{R}^{d}$ and $\forall\ell\in\{1,\dots,d\}$,
$F(W,x,a)_{\ell}:={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1},$
where $m$ is number of hidden neurons, $\exp(\cdot)$ is element-wise exponential function, $a_{\ell},W$ are the first and second layer weights respectively and $x$ is the input data. Note that, the self-attention could be written as $F(W^{K}X,W^{Q}X,W^{V}X)\in\mathbb{R}^{d\times n^{\prime}}$, where $W^{K},W^{Q},W^{V}\in\mathbb{R}^{d\times d}$ denotes key, query, and value matrix and $X\in\mathbb{R}^{d\times n^{\prime}}$ is a sequence of $n^{\prime}$ tokens. Thus, studying the two-layer softmax network is the prerequisite to understanding self-attention. See more discussion in Section H.
@@
To answer our ()(*) question above, we adopt the powerful NTK analysis paradigm in this work. Our analysis shows that, because of the normalization effect of the denominator, the Neural Tangent Kernel induced by the softmax has a good perturbation property (Lemma 5.1), which means the loss landscape of the softmax version has a large convex region. Thus, the softmax NN requires almost the same number of neurons and training steps to fit the data and converge as ReLU or exponential NN, which is illustrated in Table 1 clearly (Theorem 4.2). To demonstrate the broad applicability of our theoretical findings, we apply our analysis in a practical case study to show the generalization ability of softmax NN, where the task is learning score estimation functions in diffusion models with noisy labels, a promising approach for generative modeling, as we can smartly transfer it to a multi-dimensional regression task (Theorem 6.6). Thus, we show that gradient-based algorithms can learn the score function with a provable accuracy.
@@
  • Softmax NTK: We build up the first NTK analysis framework for two-layer NN with softmax activation function (Theorem 4.2). Furthermore, our multi-dimensional regression setting is more general than previous work [97, 48] (ReLU and $\exp$) and can be degenerated to the linear regression setting.
  • Diffusion Models Case Study: We apply our results in learning score estimation functions in diffusion models with noisy labels to verify our analysis effectiveness (Theorem 6.6).
2 Related Works
2.1 Neural Tangent Kernel
2.2 Softmax and Attention in LLMs
@@
2.3 Diffusion Model
Score-based generative diffusion models can generate high-quality image samples comparable to GANs which requires adversarial optimization [63, 133, 76]. Based on the U-Net [116], stable diffusion can successfully generate business-used images. Based on the softmax-based self-attention [107], OpenAI released a video diffusion model, SORA [102], with a surprising performance. Another line of work study training diffusion models with a better theoretical guarantee [129, 130, 132, 131, 128, 78, 77, 127, 88, 22, 27, 24, 119, 155, 14, 54, 25, 57, 150, 28]. In this work, we adapt our analysis in diffusion models.
Roadmap.
We organize our paper as follows: In Section 3, we introduce the notation system and problem setup. In Section 4, we present our main result, proving that a Softmax neural network with $\operatorname{poly}(nd)$ neurons can fit any training dataset consisting of $n$ $d$-dimensional samples for $d$-dimensional regression tasks. In Section 5, we outline the key techniques used to establish our main result. In Section 6, we extend our findings to Diffusion Models, demonstrating that Softmax neural networks can accurately learn score estimation even with noisy labels. Finally, in Section 7, we conclude the paper.
@@ 3 Preliminary
Let a vector $z\in\mathbb{R}^{n}$. We denote the $\ell_{2}$ norm as $\|z\|_{2}:=(\sum_{i=1}^{n}z_{i}^{2})^{1/2}$, the $\ell_{1}$ norm as $\|z\|_{1}:=\sum_{i=1}^{n}|z_{i}|$, $\|z\|_{0}$ as the number of non-zero entries in $z$, $\|z\|_{\infty}$ as $\max_{i\in[n]}|z_{i}|$. We use $z^{\top}$ to denote the transpose of a $z$. We use $\langle\cdot,\cdot\rangle$ to denote the inner product. Let $A\in\mathbb{R}^{n\times d}$, we use $\operatorname{vec}(A)$ to denote a length $nd$ vector. We denote the Frobenius norm as $\|A\|_{F}:=(\sum_{i\in[n],j\in[d]}A_{i,j}^{2})^{1/2}$. For a function $f(x)$, $f$ is $L$-Lipschitz if $\|f(x)-f(y)\|_{2}\leq L\cdot\|x-y\|_{2}$. Let $\cal D$} denote a distribution. We use $x\sim{\cal D}$ to denote that we sample a random variable $x$ from distribution $\cal D$}. We use $\operatorname*{{\mathbb{E}}}[]$ to denote expectation and $\Pr[]$ to denote probability. We use p.s.d. to denote the positive-semidefinite matrix.
As we have multiple indexes, to avoid confusion, we usually use $i,j\in[n]$ to index the training data, $\ell\in[d]$ to index the output dimension, $r\in[m]$ to index neuron number.
3.1 Model, Data, and Algorithm
Models. We consider a two-layer softmax neural network. The hidden layer has $m$ neurons, and we use the softmax function as the activation function, $F(W,\cdot,a):\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ and $\forall\ell\in[d_{2}]$
$F(W,x,a)_{\ell}:={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1},$ (1)
where $\exp(\cdot)$ is element-wise exponential function. We use $m$ as a normalization factor. Note that we can reduce the $d_{2}$ to $1$ for the linear regression setting. To simplify the proof, we let $d_{1}=d_{2}$. Note that our proof can generalize to different $d_{1},d_{2}$ easily.
@@ 3.1 Model, Data, and Algorithm
Data. We have $n$ training data points ${\cal D}_{n}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$. [^1] Our analysis can extend to $x_{i}\in\mathbb{R}^{d_{1}}$ and $y_{i}\in\mathbb{R}^{d_{2}}$ easily. We denote $X=[x_{1},\dots,x_{n}]\in\mathbb{R}^{d\times n}$ and $Y=[y_{1},\dots,y_{n}]\in\mathbb{R}^{d\times n}$. We assume that $\|x_{i}\|_{2}\leq 1$ and $\|y_{i}\|_{2}\leq 1$, $\forall i\in[n]$.
Gradient Descent. We use $e_{r}$ to denote a vector where the $r$-th coordinate is $1$ and everywhere else is 0. $\forall r\in[m],\forall\ell\in[d]$, we have $\frac{\partial F(W,x,a)_{\ell}}{\partial w_{r}}\in\mathbb{R}^{d}$ can be written as
$~\frac{\partial F(W,x,a)_{\ell}}{\partial w_{r}}$
$=$ $~+{{}{}{}m}\langle a_{\ell}\circ e_{r},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}x$
$~-{{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-2}$
$~\cdot\langle\exp(W^{\top}x),e_{r}\circ{\bf 1}_{m}\rangle x$
@@
We have the softmax function $\mathcal{S}\in\mathbb{R}^{m\times n}$, where $\mathcal{S}_{i}\in\mathbb{R}^{m}$ denotes $\langle\exp(W^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(W^{\top}x_{i})$ and $\mathcal{S}_{i,r}\in\mathbb{R}$ denotes $\langle\exp(W^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(w_{r}^{\top}x_{i})$, $\forall r\in[m],\forall i\in[n]$. For simplicity, we denote ${\alpha}_{i}$ as $\langle{\bf 1}_{m},\exp({W}^{\top}x_{i})\rangle$, $\exp_{i}$ as $\exp(W^{\top}x_{i})$ and $\exp_{i,r}$ as $\exp(w_{r}^{\top}x_{i})$, $\forall r\in[m],\forall i\in[n]$, when the context is clear.
We use $W(\tau)$ to denote the weights of the first layer on the timestamp $\tau$ and similar for $\mathcal{S}(\tau)$ and $F(\tau)$ when the context is clear. Now, we introduce some necessary definitions used.
We first introduce the function over the whole training dynamic.
@@
$~\Delta w_{r}(\tau)$
$:=$ $~\frac{\mathrm{d}{\cal L}(W(\tau)}{\mathrm{d}w_{r}(\tau)}$
$=$ $~{{}{}{}m}\sum_{i=1}^{n}\sum_{\ell=1}^{d}(F_{\ell,i}(\tau)-y_{\ell,i})$
$\cdot$ $~\Big(\langle a_{\ell}\circ e_{r},\mathcal{S}_{i}(\tau)\rangle-\langle a_{\ell},\mathcal{S}_{i}(\tau)\rangle\cdot\langle\mathcal{S}_{i}(\tau),e_{r}\circ{\bf 1}_{m}\rangle\Big)\cdot x_{i}$
where $\mathcal{S}_{i}(\tau)=\langle\exp(W(\tau)^{\top}x_{i}),{\bf 1}_{m}\rangle^{-1}\cdot\exp(W(\tau)^{\top}x_{i})$.
We can simplify the gradient calculation by the fact $1=\langle{\bf 1}_{m},\mathcal{S}_{i}(\tau)\rangle$. Thus, we have the following claim.
@@ 3.2 Neural Tangent Kernel
$~[H_{\ell_{1},\ell_{2}}]_{i,j}(W)$
$:=$ $~\frac{1}{m}x_{i}^{\top}x_{j}\sum_{r=1}^{m}\langle v_{\ell_{1},r},\mathcal{S}_{i}\rangle\cdot{{}{}{}m}\mathcal{S}_{i,r}\cdot\langle v_{\ell_{2},r},\mathcal{S}_{j}\rangle\cdot{{}{}{}m}\mathcal{S}_{j,r}.$
@@
Note that $H^{*}$ is a positive semi-definite matrix, and we denote its minimum eigenvalue as $\lambda:=\lambda_{\min}(H^{*})$ and we assume $\lambda>0$ as previous works [45, 3, 4].
@@ 4 Main Results
Let $C>10$ denote a sufficiently large constant. We define parameter $B$ as follows $B:=\max\{C\sigma\sqrt{\log(nd/\delta)},~1\}.$
@@
Let $\lambda=\lambda_{\min}(H^{*})>0$, $m=\Omega(\lambda^{-2}n^{2}d^{2}\exp(18B)\log^{2}(nd/\delta))$, $\eta=0.1\lambda/(mn^{2}d^{2}\exp(16B))$, and $\widehat{T}=\Omega((m\eta\lambda)^{-1}\log(nd/\epsilon))=\Omega(\lambda^{-2}n^{2}d^{2}\exp(16B)\cdot\log(nd/\epsilon))$. For any $\epsilon,\delta\in(0,0.1)$, after $\widehat{T}$ iterations, with probability at least $1-\delta$, we have $\|F(\widehat{T})-Y\|_{F}^{2}\leq\epsilon.$
@@
Consider the 1-dimension linear regression setting, i.e., $d_{1}=d$ and $d_{2}=1$. Let $\lambda=\lambda_{\min}(H^{*})>0$, $m=\Omega(\lambda^{-2}n^{2}\exp(18B)\log^{2}(n/\delta))$, $\eta=0.1\lambda/(mn^{2}\exp(16B))$, and $\widehat{T}=\Omega((m\eta\lambda)^{-1}\log(n/\epsilon))=\Omega(\lambda^{-2}n^{2}\exp(16B)$ $\cdot\log(n/\epsilon))$. For any $\epsilon,\delta\in(0,0.1)$, after $\widehat{T}$ iterations, with probability at least $1-\delta$, we have $\|F(\widehat{T})-Y\|_{2}^{2}\leq\epsilon.$